OpenStudy (vane11):

Help please? The five-number summary of scores on a test is 35 60 65 70 90 Based on this information, A. there are no outliers. B. there are low outliers. C. there are both high and low outliers.

OpenStudy (vane11):

It seems like both 35 and 90 are, but it could just be 35 since it's 5 more than the difference between 70 and 90

jimthompson5910 (jim_thompson5910):

the interquartile range (IQR) is IQR = Q3 - Q1 IQR = 70 - 60 IQR = 10

jimthompson5910 (jim_thompson5910):

anything greater than m + 1.5*IQR (where m is the median) is considered an outlier

jimthompson5910 (jim_thompson5910):

in this case m = 65 IQR = 10 so, m + 1.5*IQR = 65 + 1.5*10 = 80 because 90 is larger than 80, we know that 90 is an outlier

jimthompson5910 (jim_thompson5910):

it works the same in the other direction as well anything smaller than m - 1.5*IQR is considered an outlier

OpenStudy (vane11):

wait wouldn't Q1 equal the average of the left side of the median which is (60+35)/2 ?

jimthompson5910 (jim_thompson5910):

no, the five number summary is min, Q1, median, Q3, max so the second value is Q1

OpenStudy (vane11):

so it'd be both 90 and 35 right?

OpenStudy (vane11):

and thanks for explaining the 5 number summary :)

jimthompson5910 (jim_thompson5910):

correct, you have 2 outliers (one too low, the other too high)

jimthompson5910 (jim_thompson5910):

and yw